# Searching for gravitational waves from stellar-mass binary black holes early inspiral

### v1

@@ Abstract
The early inspiral from stellar-mass binary black holes can emit milli-Hertz gravitational wave signals, making them detectable sources for space-borne gravitational wave missions like TianQin. However, the traditional matched filtering technique poses a significant challenge for analyzing this kind of signals, as it requires an impractically high number of templates ranging from $10^{31}$ to $10^{40}$. We propose a search strategy that involves two main parts: initially, we reduce the dimensionality of the simulated signals using incremental principal component analysis (IPCA). Subsequently we train the convolutional neural networks (CNNs) based on the compressed TianQin data obtained from IPCA, aiming to develop both a detection model and a point parameter estimation model. The compression efficiency for the trained IPCA model achieves a cumulative variance ratio of 95.6% when applied to $10^{6}$ simulated signals. To evaluate the performance of CNN we generate the receiver operating characteristic curve for the detection model which is applied to the test data with varying signal-to-noise ratios. At a false alarm probability of 5% the corresponding true alarm probability for signals with a signal-to-noise ratio of 50 is 86.5%. Subsequently, we introduce the point estimation model to evaluate the value of the chirp mass of corresponding sBBH signals with an error. For signals with a signal-to-noise ratio of 50, the trained point estimation CNN model can estimate the chirp mass of most test events, with a standard deviation error of 2.48 $M_{\odot}$ and a relative error precision of 0.12.
@@ I Introduction
Due to the long duration ranging from months to years expected for sBBH signals observed by space-borne detectors, they offer the potential for high-precision estimation of physical parameters, including characteristics such as the spins of the binary and its localization[6, 7]. The information regarding the spins is particularly valuable for unraveling the evolutionary history of binary systems, given the notable variations in the effective spin predicted by various formation mechanisms[8, 9, 10].Additionally, the localization of these signals can contribute significantly to constraining the Hubble constant [11, 12, 13, 14, 15]
Indeed, the detection of GWs signal from long-lived sBBH poses a considerable challenge when employing matched filtering, especially given the need for over $10^{31}$ templates[16] . Additionally, Owen[17] et al. proposed an archival search algorithm that utilizes information from ground-based detectors to narrow the parameter space for the signal being searched in space-based detector data, allowing for more efficient data analysis. However, this method relies on the gravitational wave signal being observed by both ground-based and space-borne detectors [14]. Furthermore semi-coherent method has been explored in zero-noise scenarios [18] with the help of particle-swarm optimization, which has not provided sensitivity analysis yet. If a sBBH merger, occurring over approximately one year, can be detected by a space-borne detector, it holds the potential to offer a pre-merger warning for ground-based detectors. This necessitates the development of a swift and cost-effective pipeline for the search of sBBH signals in the mHz frequency band. Such a pipeline would serve as a crucial link between space-borne detectors and ground-based detectors.
@@
We propose to develope a search algorithm using convolutional neural network (CNN) for the search of GWs from sBBHs by TianQin. Moreover, once CNN and other neural networks complete training, their computational efficiency is very high. If we can quickly analyze the properties of a strong sBBH from a space-borne detector, it can provide early warning to other detectors. Of note, there are researches [13, 7] in which experts estimate the source parameters of an sBBH from the actual point, presuming the data will be ready from the search phase. However, without any prior knowledge from the source, the task becomes markedly difficult. Our assured pipeline can offer useful input for Bayesian inference.
@@
We have tried three different data representations for our signal: time domain, frequency domain, and time-frequency spectrum. The experiments that we have performed demonstrated that the frequency domain data was the most suitable for our purposes. We decided to work with the data compressed by IPCA model. For this reason we tested different representations of frequency domain data and we have found out that for the data processed with TianQin response it works best with the amplitude part of the data due to specific modulations in this response. Subsequent to this, we employed the trained IPCA model to compress the amplitude of detector data in the frequency domain. Upon projecting from the IPCA model, we trained a detection CNN model as well as a point parameter estimation CNN model. This serves as a demonstration of the compression- plus-search pipeline methodology for the long-lasting GW search of sBBHs.
@@ II.2 TianQin response
$\tilde{h}(f)=\sum_{l}\sum_{m}\mathcal{T}^{\rm{A,E,T}}(f,t_{lm}(f))\tilde{h}_{lm}(f).$ (7)
@@
The aforementioned extrinsic parameters are defined in the solar system barycentric coordinate system. By combining each harmonic $h_{lm}(f)$ with its corresponding transfer function $\mathcal{T}(f,t_{lm}(f))$, we can simulate the received signal samples of each TDI channel. We use IMRPhenomD to describe the original waveform (only considering the 22 mode, i.e., $l=2,m=2$ ), which is expressed as:
$\tilde{h}_{A,E,T}=\mathcal{T}^{A,E,T}(f,t_{22}(f))\tilde{h}_{22}(f).$ (9)
@@
Using a GW190521-like sBBH system as an example, as shown in Figure 1, this signal exhibits multiple cycles within the sensitive band of TianQin. The prominent oscillations are a result of detector modulations, primarily influenced by the transfer function.
@@ 13 Output $\ldots$ 2
In the task of point estimation, we employ a CNN architecture similar to that used for signal detection. However, all ‘ReLU’ activation functions are replaced with the ‘Tanh’ function, and we remove any activation function from the output layer. This is done because for point estimation we noticed that ‘ReLU’ function did not yield sufficiently good results therefore we replaced with a smooth function. The ‘Tanh’ function maps real-valued numbers to a range between -1 and 1, essentially compressing the inputs into a narrower range, which can facilitate convergence during training in our case.
@@ IV.1 Data simulation
$D_{L}$ luminosity distance to source re-scaled inversely by the SNR, Mpc
@@
Figure 2: The variance ratio of each component is arranged in descending order. The cumulative variance of $k$ components across the entire dataset amounts to 95.6%. The 100th component has a variance ratio of 0.000248, which is less than $10^{-3}$.The cumulative variance ratio of the preceding 100 components is 93.2%, indicating that the following components contribute less to the feature representation.
@@
Figure 4: A sky map showing the fitting factors for gravitational wave signals from sources distributed across the sky . In the map, the red star symbolizes the position of the DWD system J0806, which aligns with the direction of the TianQin constellation. The black solid curve delineates the orbit of TianQin. Darker blue points indicating smaller fitting factors, suggesting a lower similarity between the reconstruction and the original signal. Additionally, the size of each point corresponds to the chirp mass of the source. It’s worth noting that sources located on TianQin’s orbit tend to demonstrate lower fitting factors.
@@ IV.2 Compression
We adopt IPCA to extract the principal components from the signals, thereby reducing the dimensionality. We tested on different data representations, including time-domain signals, real and imaginary parts of frequency-domain signals, amplitude and phase of frequency-domain signals, etc. In our evaluations of various representations of responded signals, it has been observed that IPCA encounters challenges in extracting features from signals exhibiting substantial fluctuations. Consequently, we have chosen to exclusively derive the IPCA model from the amplitude of frequency-domain signals. This yields the corresponding $\rm IPCA_{i}$, where ‘i’ notes the detector output channel, designated as either A or E. We used the IPCA method from the cuML  [55] third-party package to load signals into GPU memory in batches, thereby extracting the principal components of signals from a single channel after applying the detector response. These principal components represent the main characteristics of the signal amplitude.In the Figure 5, an example is depicted within the considered frequency band. This illustration provides a comparison between the original signal amplitude and the reconstructed signal amplitude, based on estimated IPCA components, together with their residual. We can see that boundaries at the signal’s cutoff frequencies have higher residuals which means greater loss resulting from the compression process. This probably happens due to the sharp edges of the original signal.
Figure 5: The original amplitude from a random signal in ‘A’ channel with SNR 50 and the reconstruction from IPCA.The system in question has a chirp mass and symmetric mass ratio of approximately 8.3 $M_{\odot}$ and 0.18 respectively, and coalesces in a time-span of 28,923,738 seconds. Additional source parameters are referenced in Table 2, including ( $m_{1}=18.05,m_{2}=5.4\cos(\iota)=0.1,\psi=2.74,\phi_{c}=2.64,\chi_{1z}=-0.36,\chi_{2z}=0.59$, and sky position at $\lambda=2.07,\beta=-0.63$.)
Taking the extraction process of the amplitude IPCA model from the TDI A channel signal as an example, the size of the amplitude of a single signal in the A channel is (1,972001) obtained over a 3-month duration at a sampling rate of 0.25 Hz. By sampling within the astrophysical parameter range, we generate 4800000 simulated signals and a set of amplitudes, with the dimensionality is (4800000 x 1 x 972001). The variance ratio, denoted as $r_{i}$, signifies the variance $\sigma_{i}$ of the $i$-th basis in relation to the total variance $\Sigma$ of all signals, given by
$r_{i}=\frac{\sigma_{i}}{\sum}.$ (15)
@@
In our case, fitting factor is a better measure to the performance of the component extraction. We obtain the projection value of a GW signal amplitude in a channel, based on the corresponding IPCA principle components or eigenvectors. By making use of the amplitude projection values (equivalent to the coefficients of the principal components), we can reconstruct its amplitude of the original signal and then measure the similarity between the reconstruction and the original signal amplitude by calculating the fitting factor. The formula for calculating the fitting factor is shown below.
$\rm{FF}=\frac{\langle{s}(\mathbf{\theta})|s^{\prime}(\mathbf{\theta})\rangle}{\sqrt{\langle s(\mathbf{\theta})|s(\mathbf{\theta})\rangle}\sqrt{\langle s^{\prime}(\mathbf{\theta})|s^{\prime}(\mathbf{\theta})\rangle}}.$ (16)
@@
$\langle s(\mathbf{\theta})|s^{\prime}(\mathbf{\theta})\rangle=4{\sum_{f_{\rm min}}^{f_{\rm max}}\frac{|s(\mathbf{\theta})|}{\sqrt{S_{n}(f)}}\frac{|s^{\prime}(\mathbf{\theta})|}{\sqrt{S_{n}(f)}}df}.$ (17)
where $f_{\min}=0.02$ Hz and $f_{\max}=0.1$ Hz.
@@
Finally, we trained the IPCA model using 4800000 simulated signals, yielding a transformation matrix of (622080, 480). The cumulative variance ratio of these 480 components amounts to 0.958, indicating that data compressed through IPCA retains most of the variances from the amplitude of GW signals, as shown in Figure 2. Each component is a basis vector visually represented in Figure 3. From these 480 bases, it’s apparent that all are focused on the high fluctuations in relatively lower frequency band, rather than on the high frequency. Using these IPCA model, the simulated data from a single detector (2 x 9720001) can be compressed into a matrix of size (2,480). This process can also be referred to as the data pre-processing module.
To demonstrate the compression loss, we examine the fitting factor across various astrophysical parameters. Lower fitting factors predominantly correlate with high oscillations in the signal amplitude, derived from TianQin’s response. This response is represented by transfer functions that relies on 6-dimensional parameters, including the sky location. In Figure 4, the sources in specific locations (like the source localization in Figure 1 ), related to TianQin’s orbit, have a tendency to display high oscillation. This behavior subsequently leads to a lower fitting factor, exemplified by the dark blue points.
@@ IV.4 Training
Table 3: The training configuration. The total dataset will be divided, with 10% allocated for validation data and the remaining 90% used for training.
@@
best epoch / maximum epoch number 63/100 86/100
batch size 1000 1000
learning rate $\gamma=0.0001$
@@
Figure 7: The binary cross-entropy loss was calculated using fixed SNR 50 samples over 50 epochs. The best model is at 26th epoch.
Figure 8: The binary cross-entropy loss was calculated using samples with a uniformly sampled SNR ranging from 30 to 50 over 30 epochs. The best model is at 63th epoch.
@@
In the final experiment, we utilize the top-performing CNN with $10,000,000$ SNR 50 training samples as the pre-training model, employing $12,000,000$ training samples with uniformly sampled SNR ranging from 30 to 50 to continue training the model in total 300 epochs. This process yields the converged detection CNN model. The loss curves for the uniformly sampled SNR data are shown in Figure 8, The validation loss aligns closely with the training loss and keeps stable.
In task 2 (point estimation of the physical parameter ), the training strategies employed are identical to those utilized in task 1. This was decided to ensure the convolution kernels in both tasks of signal detection and point estimation would capture the same area of receptive field pertinent to sBBH GW signals. However, a smoother latent space was maintained in point estimation as compared to signal detection, due to the different activation functions in use.

### v2

@@ Abstract
The early inspiral from stellar-mass binary black hole(sBBHs) can emit milli-Hertz gravitational wave signals, making them detectable sources for space-borne gravitational wave missions like TianQin. However, the traditional matched filtering technique poses a significant challenge for analyzing this kind of signals, as it requires an impractically high number of templates ranging from $10^{31}$ to $10^{40}$. We propose a search strategy that involves two main parts: initially, we reduce the dimensionality of the simulated signals using incremental principal component analysis (IPCA). Subsequently we train the convolutional neural networks (CNNs) based on the compressed TianQin data obtained from IPCA, aiming to develop both a detection model and a point parameter estimation model. The compression efficiency for the trained IPCA model achieves a cumulative variance ratio of 95.6% when applied to $10^{6}$ simulated signals. To evaluate the performance of CNN we generate the receiver operating characteristic curve for the detection model which is applied to the test data with varying signal-to-noise ratios. At a false alarm probability of 5% the corresponding true alarm probability for signals with a signal-to-noise ratio of 50 is 86.5%. Subsequently, we introduce the point estimation model to evaluate the value of the chirp mass of corresponding sBBH signals with an error. For signals with a signal-to-noise ratio of 50, the trained point estimation CNN model can estimate the chirp mass of most test events, with a standard deviation error of 2.49 $M_{\odot}$ and a relative error precision of 0.13.
@@ I Introduction
Due to the long duration ranging from months to years expected for stellar-mass binary black hole (sBBH) signals observed by space-borne detectors, they offer the potential for high-precision estimation of physical parameters, including characteristics such as the spins of the binary and its localization[6, 7]. The information regarding the spins is particularly valuable for unraveling the evolutionary history of binary systems, given the notable variations in the effective spin predicted by various formation mechanisms[8, 9, 10].Additionally, the localization of these signals can contribute significantly to constraining the Hubble constant [11, 12, 13, 14, 15]
Indeed, the detection of GWs signal from long-lived sBBH poses a considerable challenge. It is important to note that the high computational cost required in practice restricts the applicable parameter area of the coherent search, taking into account both the large number of data points in each signal and a lot of templates [16]. Some template-based search algorithms work well when assuming a specific range of parameters [17, 18]. Owen[17] et al. proposed an archival search algorithm that utilizes information from ground-based detectors to narrow the parameter space for the signal being searched in space-based detector data, allowing for more efficient data analysis. However, this method relies on the gravitational wave signal being observed by both ground-based and space-borne detectors [14]. Furthermore, the semi-coherent method has been explored in zero-noise scenarios using particle-swarm optimization [18]. This was verified by a test for an sBBH injection with a chirp mass prior at a width of 2 solar masses. If a sBBH merger, occurring over approximately one year, can be detected by a space-borne detector, it holds the potential to offer a pre-merger warning for ground-based detectors. This necessitates the development of a swift and cost-effective pipeline for the search of sBBH signals in the mHz frequency band. Such a pipeline would serve as a crucial link between space-borne detectors and ground-based detectors.
@@
We propose to develop a search algorithm using convolutional neural network (CNN) for the search of GWs from sBBHs by TianQin. Moreover, once CNN and other neural networks complete training, their computational efficiency is very high. If we can quickly analyze the properties of a strong sBBH from a space-borne detector, it can provide early warning to other detectors. Of note, there are researches [13, 7] in which experts estimate the source parameters of an sBBH from the actual point, presuming the data will be ready from the search phase. However, without any prior knowledge from the source, the task becomes markedly difficult. Our pipeline can operate as a first stage search at a manageable cost, helping to narrow down the parameter range and provide point parameter estimates.
@@
When attempting to compress sBBH signals due to the high oscillations caused by the TianQin response, we needed to find a better signal representation to extract the dominant features. We have tried three different data representations for our signal: time domain, frequency domain (real and imagnary components), frquency domian(amplitude and phase). The amplitude is one thing we can acquire via IPCA. We undoubtedly consider the phase, but it poses a problem. The wrapped phase derived from the original responded signal also highly oscillates, and the unwrapped phase needs calibration if we consider the real data contains detector noise. In reality, noise realization will induce random artifacts for later analysis in the form of the unwrapped phase. We save the phase contribution for the future and only explore the performance with amplitude for now.
Subsequent to this, we employed the trained IPCA model to compress the amplitude of detector data in the frequency domain. Upon projecting from the IPCA model, we trained a detection CNN model as well as a point parameter estimation CNN model. This serves as a demonstration of the compression-plus-search pipeline methodology for the long-lasting GW search of sBBHs.
@@ II.2 TianQin response
$\tilde{s}(f)=\sum_{l}\sum_{m}\mathcal{T}^{\rm{A,E,T}}(f,t_{lm}(f))\tilde{h}_{lm}(f).$ (7)
@@
The aforementioned extrinsic parameters are defined in the solar system barycentric coordinate system. By combining each harmonic $\tilde{h}_{lm}(f)$ with its corresponding transfer function $\mathcal{T}(f,t_{lm}(f))$, we can simulate the received signal samples of each TDI channel. We use IMRPhenomD to describe the original waveform (only considering the 22 mode, i.e., $l=2,m=2$ ), which is expressed as:
$\tilde{s}^{\rm{A,E,T}}=\mathcal{T}^{\rm{A,E,T}}(f,t_{22}(f))\tilde{h}_{22}(f).$ (9)
@@
Using a GW190521-like sBBH system as an example, as shown in Figure 1, this signal with the optimal SNR of 50 exhibits multiple cycles within the sensitive band of TianQin. The formulation for calculating the optimal SNR is presented as follows:
$\rho=\sqrt{\langle s|s\rangle)}$ (10)
[…]
@@ 14 Output $\ldots$ 2
In the task of point estimation, we employ a CNN architecture similar to that used for signal detection. To elaborate, firstly, all ‘ReLU’ activation functions are replaced with the ‘Tanh’ function, and we remove any activation function from the output layer. This is done because for point estimation we noticed that ‘ReLU’ function did not yield sufficiently good results therefore we replaced with a smooth function. The ‘Tanh’ function maps real-valued numbers to a range between -1 and 1, essentially compressing the inputs into a narrower range, which can facilitate convergence during training in our case. Secondly, we did not use two dropout layers in the point estimation task, as this would discard too many features. This strategy allows a more sophisticated CNN model to handle a more challenging point parameter estimation task than the signal detection task.
@@ IV.1 Data simulation
$D_{L}$ luminosity distance to source re-scaled inversely by the SNR 50, Mpc
@@
Figure 2: The variance ratio of each component is arranged in descending order. The cumulative variance of $k$ components across the entire dataset amounts to 95.6% in each IPCA model. Because A and E channel signals are orthogonal, it is unsurprising that both of the extracted IPCA models have the same variance ratio for each component. The 100th component has a variance ratio of 0.000248 in $\rm{IPCA}^{A}$ model, which is less than $10^{-3}$.The cumulative variance ratio of the preceding 100 components is 93.2%, indicating that the subsequent components contribute less to the feature representation.
@@
Figure 4: The original amplitude from a random signal in ‘A’ channel with SNR 50 and the reconstruction from ${\rm IPCA}^{\rm A}$.The system in question has a chirp mass and symmetric mass ratio of approximately 8.3 $M_{\odot}$ and 0.18 respectively, and coalesces in a time-span of 28,923,738 seconds. Additional source parameters are referenced in Table 2, including ( $m_{1}=18.05\ M_{\odot}$, $m_{2}=5.4\ M_{\odot}$, $\cos(\iota)=0.1$, $\psi=2.74$ Rad, $\phi_{c}=2.64$ Rad, $\chi_{1z}=-0.36$, $\chi_{2z}=0.59$, and sky position at $\lambda=2.07$ Rad,$\beta=-0.63$ Rad.)
Figure 5: A sky map showing the amplitude fitting factors for gravitational wave signals from sources distributed across the sky . In the map, the red star symbolizes the position of the DWD system J0806, which aligns with the direction of the TianQin constellation. The black solid curve delineates the orbit of TianQin. Darker blue points indicating smaller amplitude fitting factors, suggesting a lower similarity between the reconstruction and the original signal. Additionally, the size of each point corresponds to the chirp mass of the source. It’s worth noting that sources located on TianQin’s orbit tend to demonstrate lower amplitude fitting factors.
@@ IV.2 Compression
We adopt IPCA to extract the principal components from the signals, thereby reducing the dimensionality. We tested on different data representations, including time-domain signals, real and imaginary parts of frequency-domain signals, amplitude and phase of frequency-domain signals, etc. In our evaluations of various representations of responded signals, it has been observed that IPCA encounters challenges in extracting features from signals exhibiting substantial fluctuations. Consequently, we have chosen to exclusively derive the IPCA model from the amplitude of frequency-domain signals. This yields the corresponding $\rm IPCA^{i}$, where ‘i’ notes the detector output channel, designated as either A or E. We used the IPCA method from the cuML  [55] third-party package to load signals into GPU memory in batches, thereby extracting the principal components of signals from a single channel after applying the detector response. These principal components represent the main characteristics of the signal amplitude.
Taking the extraction process of the amplitude ${\rm IPCA}^{\rm A}$ model from the TDI A channel signal as an example, the size of the amplitude of a single signal in the A channel is (1,972001) obtained over a 3-month duration at a sampling rate of 0.25 Hz. By sampling within the astrophysical parameter range in Table 2, we generate 4800000 simulated signals and a set of amplitudes, with the dimensionality is $(4800000\times 1\times 972001)$. The E channel signals will undergo a similar process for the extraction of the ${\rm IPCA}^{\rm E}$ model, maintaining an equivalent training set size as the A channel signal’s compression, represented by $(4800000\times 1\times 972001)$.
The variance ratio, denoted as $r_{j}$ in each IPCA model, signifies the variance $\sigma_{j}$ of the $j$-th basis in relation to the total variance $\Sigma$ of all signals, given by
$r_{j}=\frac{\sigma_{j}}{\sum}.$ (16)
@@
In our case, amplitude fitting factor is a better measure to the performance of the component extraction. We obtain the projection value of a GW signal amplitude in a channel, based on the corresponding IPCA principle components or eigenvectors. By making use of the amplitude projection values (equivalent to the coefficients of the principal components), we can reconstruct its amplitude of the original signal and then measure the similarity between the reconstruction and the original signal amplitude by calculating the amplitude fitting factor. The formula for calculating the amplitude fitting factor is shown below.
$\rm{FF}=\frac{\langle\left|{s}({\theta})\right||\left|s^{\prime}({\theta})\right|\rangle}{\sqrt{\langle\left|s({\theta})\right||\left|s({\theta})\right|\rangle}\sqrt{\langle\left|s^{\prime}({\theta})\right||\left|s^{\prime}({\theta})\right|\rangle}}.$ (17)
@@
$\langle\left|s({\theta})\right||\left|s^{\prime}({\theta})\right|\rangle$
$=4{\int_{f_{\rm min}}^{f_{\rm max}}\frac{\left|s({\theta})\right|\cdot\left|s^{\prime}({\theta})\right|}{S_{n}(f)}{\mathrm{d}}f}$
$\geq 4{\int_{f_{\rm min}}^{f_{\rm max}}\frac{\left|s({\theta})\right|\cdot\left|s^{\prime}({\theta})\right|\cdot\cos\Phi}{S_{n}(f)}{\rm d}f}$
$=\langle s({\theta})|s^{\prime}({\theta})\rangle,$ (18)
where $f_{\min}=0.02$ Hz and $f_{\max}=0.1$ Hz. $\Phi$ is the difference between two signals, and $-1\leq\cos\Phi\leq 1$.
@@
Finally, we trained the each IPCA model using 4800000 simulated signals, yielding a transformation matrix of (622080, 480). The training took about 57 hours using a GPU with A100 32GB memory. Employing around 1,000,000 training samples presents a reasonable compro- mise as indicated in Appendix. The cumulative variance ratio of these 480 components in ${\rm IPCA}^{\rm A}$ amounts to 0.958, indicating that data compressed through IPCA retains most of the variances from the amplitude of GW signals, as shown in Figure 2. Each component is a basis vector visually represented in Figure 3. From these 480 bases, it’s apparent that all are focused on the high fluctuations in relatively lower frequency band, rather than on the high frequency. Using these IPCA model, the simulated data from a single detector (2 x 9720001) can be compressed into a matrix of size (2, 480), achieving a compression factor of the order of 1000. This process can also be referred to as the data pre-processing module. In the Figure 4, an example is depicted within the considered frequency band. This illustration provides a comparison between the original signal amplitude and the reconstructed signal amplitude, based on estimated ${\rm IPCA}^{\rm A}$ components, together with their residual. We can see that boundaries at the signal’s cutoff frequencies have higher residuals which means greater loss resulting from the compression process. This probably happens due to the sharp edges of the original signal.
To demonstrate the compression loss, we examine the amplitude fitting factor across various astrophysical parameters. Lower amplitude fitting factors predominantly correlate with high oscillations in the signal amplitude, derived from TianQin’s response. This response is represented by transfer functions that relies on 6-dimensional parameters, including the sky location. In Figure 5, the sources in specific locations (like the source localization in Figure 1 ), related to TianQin’s orbit, have a tendency to display high oscillation. This behavior subsequently leads to a lower amplitude fitting factor, exemplified by the dark blue points.
@@ IV.4 Training
Table 3: The training configuration for the final experiment. The total dataset will be divided, with 10% allocated for validation data and the remaining 90% used for training.
@@
best epoch / maximum epoch number 63/150 149/150
batch size 5000 5000
learning rate $\gamma=1\times 10^{-5}$
@@
Figure 7: The binary cross-entropy loss was calculated on two separate datasets for the signal detection task. The small plot insert in each subfigure showcases a magnified view of changes in the loss, marked by a red rectangle. In subfigure 7, initial training was conducted over 50 epochs on a dataset with a SNR of 50, signified by dashed lines. The peak model performance, highlighted by a dotted grey vertical line, occurred at the epoch 26, serving as the pre-trained model for the succeeding dataset. In the subfigure 7. The model was fine-tuned using a second data set with SNR ranging from 30 to 50. The model’s optimal results surfaced at the 63rd epoch, out of 150, depicted by a dash-dotted grey vertical line.
Figure 8: The mean square loss was calculated on two separate datasets for the point parameter estimation task. The small plot insert in each subfigure showcases a magnified view of changes in the loss, marked by a red rectangle. In subfigure 8, the first dataset had an SNR of 50, and the optimal model was achieved at the 43rd epoch, indicated by a dotted grey vertical line. This model also served as the pre-trained model for the subsequent dataset. In subfigure 8, the second dataset, with an SNR varying between 30 and 50, yielded its best model at the 149th epoch out of a total of 150, as represented by a dash-dotted grey vertical line.
@@
In the final experiment, we utilize the top-performing CNN with $10,000,000$ SNR 50 training samples as the pre-training model , setting an initial learning rate of 1e-4. We then employed an additional $12,000,000$ training samples with uniformly sampled SNR ranging from 30 to 50, to continue training the model for next 100 epochs with an initial learning rate of 1e-5. This process yields a detection CNN model that converges without overfitting. The loss curves for the uniformly sampled SNR data are shown in Figure 7. The validation loss aligns closely with the training loss and keeps stable.
In task 2 (point estimation of the physical parameter ), the training strategies employed are identical to those utilized in task 1. This was decided to ensure the convolution kernels in both tasks of signal detection and point estimation would capture the same area of receptive field pertinent to sBBH GW signals. However, a smoother latent space was maintained in point estimation as compared to signal detection, due to the different activation functions in use. In addition, we only used the first dropout layer to avoid overfitting when training the CNN model with the fixed SNR 50 samples and then removed this dropout layer when training samples with an SNR range from 30 to 50. In order to train efficiently, we switched the learning rate decay method when training with the second dataset. In this dataset, we reduced it with a Cosine scheduler over 100 epochs. The initial learning rate was $1\times 10^{-5}$ and the minimum learning rate was $1\times 10^{-7}$. The loss curves, displayed in Figure 8 , indicate that the model has been properly trained.